OpenStudy (onthecly2):

Need some help! For the function given, state the period: f(t)=6sin(3t-(pi/6))-1 (need help getting started, I'm pretty sure I remember what to do to solve.)

OpenStudy (irishboy123):

that? \( f(t)=6 \sin(3t-(\pi/6))-1\)

OpenStudy (onthecly2):

\[f(t)=6\sin(3t-\frac{ \pi }{ 6 })-1\]

OpenStudy (onthecly2):

so yeah you have it right

OpenStudy (irishboy123):

and?

OpenStudy (onthecly2):

what exactly are you asking?

jhonyy9 (jhonyy9):

DO YOU KNOW FORMULA sin(a-b)= ?

OpenStudy (irishboy123):

\[T = \dfrac{2 \pi }{3}\] that was obvious from the start, though? right?

OpenStudy (3mar):

I think this would help you.